OpenStudy (anonymous):

Find the x- and y- intercept y=-2x+5

OpenStudy (anonymous):

x intercepts (5/2, 0) y intercepts: (0,5)

OpenStudy (anonymous):

I got x intercept (2/5,0)

OpenStudy (anonymous):

its the fraction 5 over 2 not 2 over 5

OpenStudy (anonymous):

5 is on the top, 2 is on the bottom

OpenStudy (shamim):

just put x=0 then u will get y intercept

OpenStudy (anonymous):

ya plug it in to check

OpenStudy (anonymous):

Oh yes you are right I got it now thanks

OpenStudy (anonymous):

what about this Find the x- and y- intercept x=10

OpenStudy (anonymous):

lol wait lets go back to the othe rone

OpenStudy (anonymous):

shouldnt it be like 4/2,

OpenStudy (anonymous):

plug in x and y intercepts

OpenStudy (anonymous):

to check

OpenStudy (anonymous):

Ok let check it Plug it in

OpenStudy (anonymous):

x=10? is that the full equation?

OpenStudy (anonymous):

Yes full equation I havent seen that kind of question

OpenStudy (anonymous):

does it give u the options?

OpenStudy (anonymous):

like answer choices

OpenStudy (anonymous):

my first answer was right btw i checked

OpenStudy (anonymous):

Yes the first answer was correct

OpenStudy (anonymous):

i dont think there r any intercepts for x=10

OpenStudy (anonymous):

its like a straight line

OpenStudy (anonymous):

so the x intercept must be only 10, and the other is 0 for y

OpenStudy (shamim):

y=-2x+5 if we put x=0 then we will get y intercept so y=-2*0+5 y=5

OpenStudy (anonymous):

??

OpenStudy (shamim):

so y intercept is 5

OpenStudy (anonymous):

i gave him the right answer

OpenStudy (anonymous):

i put 5

OpenStudy (anonymous):

go look up

OpenStudy (shamim):

ya